# OMGM: Orchestrate Multiple Granularities and Modalities for Efficient Multimodal Retrieval

### v1

@@ Abstract
Vision-language retrieval-augmented generation (RAG) has become an effective approach for tackling Knowledge-Based Visual Question Answering (KB-VQA), which requires external knowledge beyond the visual content presented in images. The effectiveness of Vision-language RAG systems hinges on multimodal retrieval, which is inherently challenging due to the diverse modalities and knowledge granularities in both queries and knowledge bases. Existing methods have not fully tapped into the potential interplay between these elements. We propose a multimodal RAG system featuring a coarse-to-fine, multi-step retrieval that harmonizes multiple granularities and modalities to enhance efficacy. Our system begins with a broad initial search aligning knowledge granularity for cross-modal retrieval, followed by a multimodal fusion reranking to capture the nuanced multimodal information for top entity selection. A text reranker then filters out the most relevant fine-grained section for augmented generation. Extensive experiments on the InfoSeek and Encyclopedic-VQA benchmarks show our method achieves state-of-the-art retrieval performance and highly competitive answering results, underscoring its effectiveness in advancing KB-VQA systems.
@@ 1 Introduction
Our key contributions are as follows: (i) We propose that multimodal retrieval should be tailored to the characteristics of KB-VQA, specifically in data modality and knowledge granularity. To achieve this, we introduce a coarse-to-fine multi-step retrieval strategy that progressively enhances retrieval quality by leveraging retrieval steps with varying modalities and granularities. (ii) We introduce a trainable multimodal reranker to maximize the utilization of full-modal information while minimizing inference costs by restricting the reranking scope, thereby ensuring both effectiveness and efficiency. (iii) We conducted extensive experiments on the InfoSeek and Encyclopedic-VQA (E-VQA) benchmarks, showcasing the effectiveness of the proposed method. Furthermore, a comprehensive ablation study validates the contribution of each retrieval step, offering valuable insights for designing multimodal retrieval systems in KB-VQA tasks.
@@ 4.3 Main Results
The results of our method compared with other works are presented in Tables 2 and Table 3, primarily on entity retrieval and VQA performance. Additionally, we evaluate OMGM on OK-VQA, which lacks predefined document structure, to demonstrate its generalization. The detailed analyses is presented in the Appendix D.1.
@@
4.4 Ablation Study
We conduct extensive experiments to assess the effectiveness of our framework and the design of each step, focusing on retrieval and question-answering performance. Due to space limitations, we have placed additional ablation experiments in Appendix D.4 and Appendix D.5, and the results of some ablation experiments on InfoSeek can be found in Appendix  D.6.
@@ 4.4 Ablation Study
Effect of Reranking Scope on Retrieval Performance. The reranking scope $k$ influences both the number of entities filtered during the initial search and the range of multi-modal fused features extraction during reranking. Table 7 presents the retrieval results and the average retrieval time for the first two steps with varying $k$ values. Our framework can consistently improve the retrieval capacity while increasing k from 10 to 100, accompanied with more retrieval time as well. To balance retrieval quality and efficiency, we set the reranking scope $k$ to 20 for experiments across both datasets. Additional quantitative experimental results and analysis of the efficiency of OMGM’s step-by-step framework can be found in the Appendix D.2.
@@ Appendix D More Experiments and Ablation studys
D.1 Generalization on Additional Benchmark with Differen Document Structures
For E-VQA and InfoSeek documents, which consist of entity-based Wikipedia articles, we segmented them into sections based on their inherent structure. This method allowed us to leverage the natural organization of Wiki-style content, which often includes headings and subheadings, to create meaningful sections for retrieval. For KB-VQA datasets with document structures that differ significantly from E-VQA or InfoSeek, alternative segmentation strategies can be employed to ensure effective organization of knowledge. One potential approach is to use rule-based methods, which rely on predefined heuristics such as paragraph breaks, headings, or specific keywords to define section boundaries. Another approach is semantic clustering, which groups text segments based on semantic similarity, enabling the creation of fine-grained, section-like knowledge units even in the absence of explicit document structure.
To evaluate the generalizability of OMGM on different segmenation setting, we tested our method on the OK-VQAMarino et al. (2019) dataset, which lack a highly structured format and segmented using a rule-based segmentation method. As shown in Table 8, OMGM demonstrated strong performance, achieving higher retrieval accuracy (Pseudo Recall@5) and VQA scores compared to the strong baseline PreFLMRLin et al. (2024). Specifically, OMGM achieved a Pseudo Recall@5 of 73.4 and a VQA score of 66.57, outperforming PreFLMR by a notable margin. These results suggest that OMGM’s framework is robust and adaptable to varying document structures across different KB-VQA datasets.
Method Pseudo Recall@5 VQA score
PreFLMR 70.9 61.88
OMGM 73.4 66.57
D.2 Quantitative evaluation on the efficiency of OMGM’s step-by-step approach.
To address concerns regarding the efficiency of the proposed OMGM framework, we conducted a comparative evaluation of OMGM, the one-step multimodal RAG method PreFLMRLin et al. (2024), and the direct use of LLaVA-1.5-7B for VQA on the E-VQA dataset. We evaluated three key metrics: average retrieval time, average inference time, and VQA performance. It is worth noting that PreFLMR preprocesses and encapsulates all passage embeddings prior to inference, which reduces retrieval time during runtime to primarily consist of query embedding and similarity matching. To ensure a fair comparison, OMGM’s retrieval time calculation also focuses on these two components.
Method Avg. Ret. Time Avg. Inf. Time VQA Result
LLaVA-1.5-7B - 1.432 17.00
PreFLMR 0.984 2.196 54.45
OMGM 0.402 2.023 63.39
D.3 Consistency of our work across LLMs and MLLMs.
@@ D.1 Generalization on Additional Benchmark with Differen Document Structures
D.4 Effect of different hyper-parameters’ value in step 2 and 3
@@
D.5 Methods for computing multimodal fusion feature similarities
@@
D.6 Some Ablation study results in InfoSeek

### v2

@@ Abstract
Vision-language retrieval-augmented generation (RAG) has become an effective approach for tackling Knowledge-Based Visual Question Answering (KB-VQA), which requires external knowledge beyond the visual content presented in images. The effectiveness of Vision-language RAG systems hinges on multimodal retrieval, which is inherently challenging due to the diverse modalities and knowledge granularities in both queries and knowledge bases. Existing methods have not fully tapped into the potential interplay between these elements. We propose a multimodal RAG system featuring a coarse-to-fine, multi-step retrieval that harmonizes multiple granularities and modalities to enhance efficacy. Our system begins with a broad initial search aligning knowledge granularity for cross-modal retrieval, followed by a multimodal fusion reranking to capture the nuanced multimodal information for top entity selection. A text reranker then filters out the most relevant fine-grained section for augmented generation. Extensive experiments on the InfoSeek and Encyclopedic-VQA benchmarks show our method achieves state-of-the-art retrieval performance and highly competitive answering results, underscoring its effectiveness in advancing KB-VQA systems. Our code can be found at https://github.com/ChaoLinAViy/OMGM.
@@ 1 Introduction
In summary, the main contributions of our work can be summarized as follows:
  • We propose that multimodal retrieval should be tailored to the characteristics of KB-VQA, specifically in data modality and knowledge granularity. To achieve this, we introduce a coarse-to-fine multi-step retrieval strategy that progressively enhances retrieval quality by leveraging retrieval steps with varying modalities and granularities.
  • We introduce a trainable multimodal reranker to maximize the utilization of full-modal information while minimizing inference costs by restricting the reranking scope, thereby ensuring both effectiveness and efficiency.
  • We conducted extensive experiments on the InfoSeek and Encyclopedic-VQA (E-VQA) benchmarks, showcasing the effectiveness of the proposed method. Furthermore, a comprehensive ablation study validates the contribution of each retrieval step, offering valuable insights for designing multimodal retrieval systems in KB-VQA tasks.
@@ 4.3 Main Results
The results of our method compared with other works are presented in Tables 2 and Table 3, primarily on entity retrieval and VQA performance.
@@
4.4 Ablation Study
We conduct extensive experiments to assess the effectiveness of our framework and the design of each step, focusing on retrieval and question-answering performance. Due to space limitations, we have placed additional ablation experiments in Appendix D.2 and Appendix D.3, and the results of some ablation experiments on InfoSeek can be found in Appendix  D.4.
@@ 4.4 Ablation Study
Effect of Reranking Scope on Retrieval Performance. The reranking scope $k$ influences both the number of entities filtered during the initial search and the range of multi-modal fused features extraction during reranking. Table 7 presents the retrieval results and the average retrieval time for the first two steps with varying $k$ values. Our framework can consistently improve the retrieval capacity while increasing k from 10 to 100, accompanied with more retrieval time as well. To balance retrieval quality and efficiency, we set the reranking scope $k$ to 20 for experiments across both datasets.
@@
Generalization on Additional Benchmark with Differen Document Structures. For E-VQA and InfoSeek documents, which consist of entity-based Wikipedia articles, we segmented them into sections based on their inherent structure. This method allowed us to leverage the natural organization of Wiki-style content, which often includes headings and subheadings, to create meaningful sections for retrieval. For KB-VQA datasets with document structures that differ significantly from E-VQA or InfoSeek, alternative segmentation strategies can be employed to ensure effective organization of knowledge. One potential approach is to use rule-based method, which relies on predefined heuristics such as paragraph breaks, headings, or specific keywords to define section boundaries. Another approach is semantic clustering, which groups text segments based on semantic similarity, enabling the creation of fine-grained, section-like knowledge units even in the absence of explicit document structure.
To evaluate the generalizability of OMGM on different segmenation setting, we tested our method on the OK-VQAMarino et al. (2019) dataset, which lacks a highly structured format and segmented using a rule-based segmentation method. As shown in Table 8, OMGM demonstrated strong performance, achieving higher retrieval accuracy (Pseudo Recall@5) and VQA scores compared to the strong baseline PreFLMRLin et al. (2024). Specifically, OMGM achieved a Pseudo Recall@5 of 73.4 and a VQA score of 66.57, outperforming PreFLMR by a notable margin. These results suggest that OMGM’s framework is robust and adaptable to varying document structures across different KB-VQA datasets.
Method Pseudo Recall@5 VQA score
PreFLMR 70.9 61.88
OMGM 73.4 66.57
Quantitative evaluation on the efficiency of OMGM’s step-by-step approach. To further perform the efficiency of the proposed OMGM framework, we conducted a comparative evaluation of OMGM, the one-step multimodal RAG method PreFLMRLin et al. (2024), and the direct use of LLaVA-1.5-7B for VQA on the E-VQA dataset. We evaluated three key metrics: average retrieval time, average inference time, and VQA performance. It is worth noting that PreFLMR preprocesses and encapsulates all passage embeddings prior to inference, which reduces retrieval time during runtime to primarily consist of query embedding and similarity matching. To ensure a fair comparison, OMGM’s retrieval time calculation also focuses on these two components.
Method Avg. Ret. Time Avg. Inf. Time VQA Result
LLaVA-1.5-7B - 1.432 17.00
PreFLMR 0.984 2.196 54.45
OMGM 0.402 2.023 63.39
@@ Appendix D More Experiments and Ablation studys
D.1 Consistency of our work across LLMs and MLLMs.
@@
D.2 Effect of different hyper-parameters’ value in step 2 and 3
@@
D.3 Methods for computing multimodal fusion feature similarities
@@
D.4 Some Ablation study results in InfoSeek